OpenStudy (anonymous):

solve log(base 3)3x + log(base 3) 1 = log(base 3) (6x + 9)

OpenStudy (anonymous):

Remember that addition of logs just multiplies them, so log_3 ((3x)*(1)) = log_3 (6x + 9) 3x = 6x + 9 -3 = 3x -1 = x